OpenStudy (anonymous):

e^(x)>10 help please

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

the answer is: \[\sum_{0}^{infinity and beyond}\]

OpenStudy (anonymous):

do the logarithm dance

OpenStudy (anonymous):

take ln to both sides: \[\LARGE \ln(e^x)>\ln(10)\] \[\LARGE x>\ln(10)\]

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

medal goes to you